OpenStudy (anonymous):

What is the GCF of 40xy and 24x? A. 8x B. 4xy C. 24x D. 40xy

OpenStudy (anonymous):

C.

OpenStudy (across):

@SuperhotMilf, howcome?

OpenStudy (across):

\[40xy=2^3\cdot5\cdot x\cdot y\]\[24x=2^3\cdot3\cdot x\]\[\gcd(40xy,24x)=2^3\cdot x=8x\]